Exercise 14-16

On January 1, 2020, Blue Sky Company makes the two following acquisitions.

1. Purchases land having a fair value of $360,000 by issuing a 5-year, zero-interest-bearing promissory note in the face amount of $606,621.

2. Purchases equipment by issuing a 7%, 8-year promissory note having a maturity value of $560,000 (interest payable annually).

The company has to pay 11% interest for funds from its bank.

(a) Record the two journal entries that should be recorded by Blue Sky Company for the two purchases on January 1, 2020.

(b) Record the interest at the end of the first year on both notes using the effective-interest method.

(Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answer to 0 decimal places e.g. 58,971. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1

Ans.

a)

Date Account Titles and Explanation Debit Credit
January 1, 2020 Land $360,000.00
Discount on notes payable $246,621.00
                 Notes payable $   606,621.00
(To record purchase of land by issuing note payable)

PV of $606,621 discounted at 11% =606,621/(1.11)^5 = $ 360,000

2.

Computation of the discount on notes payable:

Maturity value $560,000

Present value of $560,000 due in 8 years at 11% = $560,000 * 0.43393 = $ 243,000

Present value of $39,200 payable annually for 8 years at 11% annually—$39,200 * 5.14612 = $ 201,728

Present value of the note = $ 243,000 + $ 201,728 = $ 444,728

Discount = $ 560,000 - $ 444,728 = $ 115,272

Date Account Titles and Explanation Debit Credit
  January 1, 2020 Equipment $444,728.00
Discount on notes payable $115,272.00
           Notes payable $   560,000.00
(To record purchase of equipment by issuing note payable)

b)

1.

Date Account Titles and Explanation Debit Credit
December 31, 2020 Interest expense ($ 360,000*11%) $39,600
Discount on notes payable $39,600
(To record the interest expense recorded and discount amortized)

2.

Date Account Titles and Explanation Debit Credit
December 31, 2020 Interest expense ($444,728 * 11%) $48,920
         Discount on notes payable $9,720
         Interest Payable ( $ 560,000 * 7%) $39,200
(To record the interest expense recorded)